# RLGNet: Repeating-Local-Global History Network for Temporal Knowledge Graph Reasoning

### v1

@@ 5.4 Hyperparameter Analysis
Effect of $\omega$ and $tok_{k}$ Values in RLGNet. Figure 4 shows the impact of $\omega$ and $top_{k}$ on the MRR results in the ICEWS14 and YAGO datasets. As $\omega$ increases, the performance of the model on ICEWS14 fluctuates slightly, while the performance on the YAGO dataset rapidly declines, so we set $\omega$ to 1 in all datasets. In the YAGO dataset, setting $top_{k}$ at a higher value yields better performance however, changes in $top_{k}$ do not significantly affect the performance on the ICEWS14 dataset. Therefore, we set $top_{k}$ for all datasets to 20.
@@ 6 Conclusion
In this paper, we propose RLGNet for TKG reasoning. RLGNet captures the characteristics of corresponding facts by using repeating, local, and global historical encoders respectively. In the future, finding a superior strategy to balance the contributions of global and local historical information is a worthwhile research problem. Experiments on six benchmark datasets prove that RLGNet outperforms existing models in most extrapolation tasks.
7 Limitations
In this paper, we have proposed a TKG prediction model based on three encoders. However, our approach has certain limitations:
  • 1.
    Due to the constraints of computational resources, the three encoders need to train independently. This method prevents them from sharing weights among themselves, potentially affecting the predictive performance of the model.
  • 2.
    In multi-step and single-step reasoning tasks, the contributions of global and local historical information should indeed be different. We have adopted a weighted average strategy to alleviate this issue to some extent, but choosing a specific hyperparameter $\alpha$ cannot achieve optimal performance on both multi-step and single-step reasoning tasks at the same time.
Acknowledgments

### v2

@@ 5.5 Effectiveness of Ensemble Learning Strategy
Model ICEWS14 YAGO
Ensem Joint Ensem Joint
Glo+Loc 45.56 44.98 87.38 87.38
Loc+Rep 45.91 45.10 89.65 88.95
Glo+Rep 43.74 42.77 78.85 82.52
Glo+Loc+Rep 46.15 45.40 89.69 88.55
Table 6: The MRR (in percentage) results of the ablation studies. Ensem and Joint represent ensemble learning and joint learning, respectively.
@@ 6 Conclusion
In this paper, we propose RLGNet for TKG reasoning. RLGNet is a model based on an ensemble learning strategy with a hybrid multi-scale architecture. By using Repeating, Local, and Global modules, RLGNet can capture historical information at different scales. To achieve complementary capabilities between modules and improve the model’s performance in single-hop and multi-hop reasoning tasks, RLGNet adopts a hybrid architecture design: the Local and Global modules utilize RNN and MLP architectures, respectively, and their scores are combined through weighted summation. We also employ an ensemble learning strategy to further enhance the model’s performance and generalization ability. Experiments on six benchmark datasets demonstrate that RLGNet outperforms existing models in most extrapolation tasks.
7 Acknowledgments